A satisfactory day at Grosmont today with 9 of us there for most of the day.
The RH connecting rod was removed from the Q6 and the clearances of both little end pins checked.
On the K1 we continued changing the defective roof stay nuts (only 7 left to do now) serviced the LH drain cocks and both rear cylinder relief valves (the left leading has also been started) refitted the remetalled LH cross head, die pen tested the stuffing box end of the regulator rod (found to be ok) progressed the annual maintenance of both clack boxes including recutting both steam valve seats, adjusted the ashpan rear hopper door linkage, prepared the back of the tender tank for seal welding (where minor leaks have occurred to previously repaired areas) and measured the RH slide bar in connection with the ongoing repairs to the RH cross head (some minor fracturing has been welded and the top slipper, which is very worn and fractured, is being replaced).
The next working party will be on Wednesday then both days next weekend. We need to progress the work on both the Q6 and K1 over the next few weeks so please come and join us if you can.